OpenStudy (jetman24):

Simplify the following expression by adding and subtracting: -5√50 + 3√2 - 8√18 A. -181√2 B. -46√2 C. -44√2 D. -18√2

OpenStudy (jetman24):

will give medal for help

OpenStudy (supersmart1001):

ok i will see what i can do

OpenStudy (darv):

Answer is B

OpenStudy (supersmart1001):

just go to answers.com and see if this is there if then u got ur answer

OpenStudy (darv):

Do you need the working?

OpenStudy (darv):

-5√50+3√2-8√18 =-5(5)√2+3√2-8(3)√2 =-25√2+3√2-24√2 =(-25+3-24)√2 =-46√2

OpenStudy (jetman24):

yes please @Darv

OpenStudy (jetman24):

thank you

OpenStudy (darv):

No problem bud